OpenStudy (anonymous):

HEEEEEELLLP FOR A MEDAL!!!!!!!!!!!!!!!!! A family has two cars. The first car has a fuel efficiency of 30 miles per gallon of gas and the second has a fuel efficiency of 35 miles per gallon of gas. During one particular week, the two cars went a combined total of 1500 miles, for a total gas consumption of 45 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

i help u n u help me deal?

OpenStudy (homeworksucks):

One of you is doing systems of linear equations and the other Bernoulli Principles (calculus maybe?).. This might not work out

OpenStudy (anonymous):

yeah its gonna work

OpenStudy (anonymous):

30x + 35y = 1500 x + y = 45 --> x = 45 - y now sub 45 - y in for x in the other equation 30(45 - y) + 35y = 1500 1350 - 30y + 35y = 1500 5y = 1500 - 1350 5y = 150 y = 150/5 y = 30 now sub 30 in for y x + y = 45 x + 30 = 45 x = 45 - 30 x = 15 check... 30(15) + 35(30) = 1500 450 + 1050 = 1500 1500 = 1500 (correct) first car (x) used 15 gallons second car (y) used 30 gallons